OpenStudy (anonymous):

To be solved 2x^2-5x+3≤ 0

OpenStudy (phi):

first step: can you factor this quadratic ?

OpenStudy (anonymous):

To solve it you can complete the square: \[2x^2-5x+3\le0\] First divide everything by 2: \[x^2-\frac{ 5 }{ 2 }x+\frac{ 3 }{ 2 }\le0\] Subtract 3/2 from both sides: \[x^2-\frac{ 5 }{ 2 }x \le-\frac{ 3 }{ 2 }\] Take half of the b-term "-5/2" and square it so half of it is -5/4 and if you square that you get 25/16. Add that to both sides: \[x^2-\frac{ 5 }{ 2 }x+\frac{ 25 }{ 16 }\le -\frac{ 3 }{ 2 }+\frac{ 25 }{ 16 }\] Simplify: \[x^2-\frac{ 5 }{ 2 }x+\frac{ 25 }{ 16 }\le \frac{ 1 }{ 16 }\] factor the right side to the half the b-term or -5/4 you found earlier: \[(x-\frac{ 5 }{ 4 })(x-\frac{ 5 }{ 4 })\le \frac{ 1 }{ 16}\] becomes: \[(x-\frac{ 5 }{ 4 })^2\le \frac{ 1 }{ 16 }\] Take the square root of both sides remember square roots are positive and negative: \[x-\frac{ 5 }{ 4 }\le \pm \frac{ 1 }{ 4 }\] Add 5/4 to both sides: \[x \le \frac{ 3 }{ 2 }\] or \[x \le1\]

OpenStudy (anonymous):

thanks

OpenStudy (phi):

to solve \[ 2x^2-5x+3≤ 0 \] first factor it into \[ (2x -3)(x-1) ≤ 0 \] Now notice that the two terms on the left will be negative (i.e. less than zero) when one term is negative and the other is positive: so you need either 2x-3≤ 0 and x-1 ≥ 0 --> x≤ 3/2 and x ≥ 1 or 2x-3 ≥ 0 and x-1 ≤ 0 --> x ≥ 3/2 and x ≤ 1 *** this case is not possible** the answer is 1 ≤ x ≤ 3/2

OpenStudy (anonymous):

@phi got the same answer i got but wrote it as a compound inequality!